OpenStudy (anonymous):

which is equivalent to √6m (multiplied by) √3m ...3√2m ...9m ...m√18 ...3m√2

OpenStudy (anonymous):

if it is \[\sqrt{6m}\times \sqrt{3m}\] then the answer is \[3m\sqrt{2}\]